OpenStudy (adelyn_helps):

Can someone help me with 1 question please and I dont just want yall to give me the answer please explain! Thx and I will give medals

OpenStudy (adelyn_helps):

The cone and cylinder above have the same radius, r, and height, h. The volume of the cone is 72 cubic centimeters. What is the volume of the cylinder? A. 144 cubic centimeters B. 24 cubic centimeters C. 288 cubic centimeters D. 216 cubic centimeters

OpenStudy (adelyn_helps):

(and let me get the picture)

OpenStudy (amorfide):

volume of cylinder \[V=\pi(r^{2})h\] volume of cone \[V=\pi(r^{2})\frac{ h }{ 3 }\] volume of cone is 72 so \[72=\pi(r^{2})\frac{ h }{ 3 }\] infact you can already see that the only difference between volume is that one is divided by 3 so if we focus on volume of cone if we multiply both sides by 3 \[72 \times 3 = \pi(r^{2})h\] \[216=\pi(r^{2})h\] so now since the volume of a cylinder is \[\pi(r^{2})h\] we now know that the volume of the cylinder is 216

OpenStudy (adelyn_helps):

Thank you so much! God Bless and have a good day!